A bag contains 30 coins, all of which are either quarters or dimes. If the total value of the coins is $5.85, how many of each kind are there?
Found 2 solutions by macston, ikleyn:
Answer by macston(5194)

.
Q=number of quarters; D=number of dimes
.
Q+D=30
D=30-Q . Use this to substitute for D below
.
$0.10D+$0.25Q=$5.85 . Substitute for D from above.
$0.10(30-Q)+$0.25Q=$5.85
$3.00-$0.10Q+$0.25Q=$5.85
$0.15Q=$2.85
Q=19
.
ANSWER 1: There were 19 quarters.
.
D=30-Q=30-19=11
.
ANSWER 2: There were 11 dimes.
.
CHECK:
.
$0.10D+$0.25Q=$5.85
$0.10(11)+$0.25(19)=$5.85
$1.10+$4.75=$5.85
$5.85=$5.85
.

Answer by ikleyn(52788)

.
For coin problems and their detailed solutions see the lessons in this site:
    - Coin problems
    - More Coin problems
    - Solving coin problems without using equations
    - Kevin and Randy Muise have a jar containing coins
    - Typical coin problems from the archive

Read them attentively and become an expert in this field.